Project

General

Profile

12052013

Nightly Observing Plan for MJD 56632 (starting evening of Dec 5, 2013)

Observers

Observer 1 Observer 2 Run Manager Assistant Run Manager SISPI Expert on Call
Rutu Das Elisabeth Krause Lyndsay Old Marcelle Soares-Santos
ext. 521 ext. 515 ext. 516 ext. 518

CTIO staff (Telescope Operators)

Day Support Night Assistant
Manuel Hernandez Claudio Aguilera

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56632, date 20131205, Site CTIO, UT-offset 3 hr
Current local time 00:54 CLST     UT 03:54     SidTim 04:11
Sunset     20:42 CLST  23:42 UT   SidTim  360 deg  
Eve. -10   21:22 CLST  00:22 UT   SidTim   10 deg  Zeropoint; standards
Eve. -14   21:45 CLST  00:45 UT   SidTim   15 deg  Begin obstac
Eve. -18   22:09 CLST  01:09 UT   SidTim   21 deg  Last chance standards
Moonset    23:11 CLST  02:11 UT   SidTim   37 deg  
Midpoint   01:34 CLST  04:34 UT   SidTim   72 deg  
Morn -18   04:59 CLST  07:59 UT   SidTim  124 deg  Begin astronomical twilight
Morn Std   05:23 CLST  08:23 UT   SidTim  130 deg  End obstac; begin standards
Morn -10   05:45 CLST  08:45 UT   SidTim  136 deg  
Sunrise    06:26 CLST  09:26 UT   SidTim  146 deg  
Moon position (deg): RA 298.9, Dec -14.8 Phase: 0.240

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards (preceed with zeros script: script to use: flush_Vsub.json).
  3. Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
  4. ObsTac.
  5. End of night standards.
  6. End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json

Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.

UT Date = '2013/12/06 00:22' Local time = 21:22
E8-A ra 20:07:22 dec -44:37:01 secz 1.7 script to use: DES_std_e8a_Yzirgu.json
SDSSJ2140-0000 ra 21:40:00 dec +00:00:00 secz 1.6 script to use: DES_std_sdssj2140_Yzirgu.json
SDSSJ0100-0000 ra 01:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0100_Yzirgu.json

End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/12/06 08:23' Local time = 05:23
SDSSJ1227-0000 ra 12:27:00 dec +00:00:00 secz 2.1 script to use: DES_std_sdssj1227_ugrizY.json
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.4 script to use: DES_std_sdssj1048_ugrizY.json
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_ugrizY.json

Notes from 4pm Meeting

Run manager's Meeting 12/05/13
-------------------------------
HTD, Klaus, Michael, Margaret, Eric
Lyndsay, Marcelle, Elisabeth K., Alex, Josh
Manuel, Robert, Steve

Telescope and Camera
--------------------
Mirror is being cleaned and they will finish soon. A big problem with
the Dome encoders that cost us images two nights ago is fixed.
The new dome encoders are being installed. LN2 lines pumped today.
Dome new environmental controls system is going-in.

SISPI is restarted. Telescope is ready to go at the white spot.

Last Night's Data
-----------------
Guider noise seen early last night; fixed by stop/reconfiguring/start pan-g.
Cost 20 minutes or so.

Obstac most of the night. Some SN fields.

Some occluded exposures: procedure is to "do not process".
DESDM will mark them as bad and they will be so in the exposure table.

DESDM
-----
12/2 and 12/3 DQ will come out shortly. Last night's data is being processed already.

Steve and Robert agree that the new DQ is ready to go after one more cloud tweak.
Robert wants to rerun on the full dataset. Starting Monday we will
begin to implement the new DQ. SN DQ will also come into this as
"first cut eval" but for the moment "first cut eval v3" in Robert's Table.
For the moment it's still a CSV to the mountain. Implement a python
code starting later next week.

When Robert gets the new evaluation for the previous data he sends it to
Eric so we can make an evaluation about whether we want to apply it
retroactively.

For Tonight
-----------
weather is looking OK - some chance of light clouds.
Obstac is expected to go to tiling 4. Some SN fields will hit if the seeing is
poor.